OpenStudy (anonymous):

solve cotX^2+4cotX-7=0

OpenStudy (anonymous):

find the general solution

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (solomonzelman):

you can at first say, let \(\cot(x)=a\), which will give you the following quadratic equation, \(a^2+4a-7=0\)

OpenStudy (solomonzelman):

Solve this equation for a, and after you find the solutions for a, you can find the value of x, using the fact that \(a=\cot(x)\)

OpenStudy (solomonzelman):

(well, not the value of x, but a set of solutions, but I hope you see what I mean)

OpenStudy (anonymous):

what about the general solution?

OpenStudy (mathstudent55):

When you write \(\cot x^2\), do you really mean \(\cot^2 x\) ?

OpenStudy (anonymous):

the latter

OpenStudy (solomonzelman):

yes, mathstudent55 I would assume so:) I didn't even notice that mistake (if it is a mistake)

OpenStudy (solomonzelman):

have you solved for "a" (using a substitution of a=cot(x) that I offered) ?

OpenStudy (anonymous):

yes but that just gave me -2 +or-sqrt(44)

OpenStudy (solomonzelman):

i got mixed up for some reason

OpenStudy (solomonzelman):

a²-4a-7=0 a²-4a=7 a²-4a+4=7+4 a²-4a+4=11 (a-2)²=11 a-2=±√11 a=-2±√11 that is what I get

OpenStudy (anonymous):

its +4a

OpenStudy (solomonzelman):

oh, tnx for catching me

OpenStudy (anonymous):

(y)

OpenStudy (solomonzelman):

oh, my fault, i did the step incorrectly

OpenStudy (solomonzelman):

a²-4a-7=0 a²-4a=7 a²-4a+4=7+4 a²-4a+4=11 (a-2)²=11 a-2=±√11 a=2±√11 that is what I get if -4a ------------------ thus with +4a i would get now with +4a i get a²+4a-7=0 a²+4a=7 a²+4a+4=7+4 a²+4a+4=11 (a+2)²=11 a+2=±√11 a=-2±√11

OpenStudy (solomonzelman):

and by the quadratic formula a²+4a-7=0 \(\large\color{black}{ \displaystyle \frac{-4\pm\sqrt{(4)^2-4(-7)(1)} }{2} }\) \(\large\color{black}{ \displaystyle \frac{-4\pm\sqrt{44} }{2} }\) \(\large\color{black}{ \displaystyle \frac{-4}{2}\pm\frac{\sqrt{44} }{2} }\)

OpenStudy (solomonzelman):

see I am getting the same thing a = -2±√11

OpenStudy (anonymous):

okay, yay thanks, i get it now!

OpenStudy (solomonzelman):

yes, so \(\cot(x)=-2+\sqrt{11}\) or \(\cot(x)=-2-\sqrt{11}\)

OpenStudy (solomonzelman):

\(\cot^{-1}(-2+\sqrt{11})=a\) or \(\cot^{-1}(-2-\sqrt{11})=a\)